A chemist wishes to make 2 L of an 8% acid solution by mixing a 10% acid solution and a 5% acid solution. How many litters of each solution should the chemist use?

Answer by nerdybill(7384)

A chemist wishes to make 2 L of an 8% acid solution by mixing a 10% acid solution and a 5% acid solution. How many litters of each solution should the chemist use?
Let x = amount (liters) of 10% solution
then
2-x = amount (liters) of 5% solution
.
.10x + .05(2-x) = .08(2)
.10x + .10 - .05x = .16
.05x + .10 = .16
.05x = .06
x = 1.2 liters (10% solution)
.
5% solution:
2-x = 2-1.2 = .8 liters